OpenStudy (anonymous):

Find the next three terms of the sequence: -3, 12, -48, 192, …. A.-768, 3072, -12288 B.768, -3072, 12288 C.576, -1728, 5184 D.-576, 1728, -5184

OpenStudy (anonymous):

Each number appears to be multiplied by -4. So then you would get: -768, 3072, -12288 or A.

OpenStudy (anonymous):

3 3*4 = 12 12*4 = 48 48*4 = 192 Continue that pattern alternating negative signs.

OpenStudy (anonymous):

first of all this called G.S;where is diffrence is calculated by dividing or multiplying ,so u should divide the second term over the 1st term so that u know the difrence = ??? so if u get 12/-3=-4 therefore the r=(-4) no go to the last term and multiply -4 to it : 192*-4=-768.,..... and so on :) ______________________- NEED MORE EXPLANATION ??!